# Cupid: Pose-Grounded Generative 3D Reconstruction from a Single Image

### v1

@@ 6 Discussion
We have presented Cupid, a pose-grounded generative framework that addresses single-image 3D reconstruction from a new perspective. By explicitly anchoring generation to a canonical 3D object representation, Cupid naturally integrates 3D priors learned in the canonical space with image inputs and enforces both geometric and photometric fidelity in generation via camera pose-aligned image conditioning. This simple but powerful grounding jointly enables high-fidelity single-view reconstruction and scene composition, significantly eases the acquisition of 3D assets.
Despite its strong performance and consistent gains over prior methods, Cupid has limitations. First, like most existing 3D object generation methods, it relies on an accurate object mask, and boundary errors in real images can degrade reconstruction quality. Second, lighting can be partially baked into textures. Further improvements require better material–light disentanglement. Third, our synthetic training images are mostly centered, making off-centered objects more challenging in real scene generation. Nevertheless, these are not fundamental limitations of our method and can be alleviated with better training data and supervision.
Looking forward, this generation-based reconstruction framework opens promising test-time extensions, particularly for multi-view scenarios. By leveraging multiple 2D images, our method can refine 3D reconstructions to align with all observations by fusing shared object volumes during sampling, similar to Multi-Diffusion (Bar-Tal et al., 2023) in image domain. As shown in Figure 6, this yields an SfM-like system, though challenges like misaligned 3D object orientations (Lu et al., 2025) from diverse views require advanced fusion schemes in future work. Furthermore, Cupid’s joint modeling of 3D structure and pose enables bidirectional capabilities, such as flexible generation with known poses, or pose estimation from images of given 3D objects, facilitating tasks in virtual and augmented reality and embodied AI for robotics.
Ethics statement
Our work advances 3D reconstruction for embodied AI, robotics, and augmented reality. However, it also poses ethical risks if misused. A primary concern is that bad actors might generate unauthorized 3D assets from 2D images, which could lead to intellectual property infringement, unauthorized replication of cultural artifacts, or privacy violations, such as non-consensual modeling.
To address these issues, our open-source code includes ethical guidelines. Additionally, we recognize that training datasets may contain biases, such as cultural underrepresentation; future efforts should prioritize inclusive data collection. We are also aware of the environmental impact of energy use during training and provide transparency regarding this matter. We are committed to evaluating societal impacts and promoting responsible usage of our technology to maximize its benefits.
Reproducibility Statement
The datasets used, including ABO (Collins et al., 2022), HSSD (Khanna et al., 2024), 3D-FUTURE (Fu et al., 2021), and Objaverse-XL (Deitke et al., 2023), are publicly available. TRELLIS, upon which our work is built, is also publicly accessible via its official GitHub repository (Xiang et al., 2025). The Blender script for synthesizing 2D images can be adapted from the TRELLIS code base. We provide detailed training implementation in Appendix A.1. Our occlusion-aware model extends Amodal3R (Wu et al., 2025), which is publicly available; our customized modifications are detailed in Appendix A.2. While our least-squares pose solver is implemented in CUDA, an equally effective alternative can be found using the calibrateCamera function in OpenCV. Finally, we will open-source all our training and inference code to facilitate reproducibility.
@@ Appendix A Appendix
Figure 7: Diversity of our Generative Reconstruction. We visualize canonical front and back views of generated 3D objects using random seeds (1-4). Given a single input image, our model synthesizes diverse hypotheses for unobserved regions while remaining highly consistent with visible regions. In contrast, the base 3D generator (Xiang et al., 2025) struggles to conform to the input image, or produce less diverse unobserved regions.
@@ A.1 Implementation
We train our model using several datasets, including ABO (Collins et al., 2022), HSSD (Khanna et al., 2024), 3D-FUTURE (Fu et al., 2021), and a subset of Objaverse-XL (Deitke et al., 2023), totaling approximately 260K 3D assets. These artist-curated datasets are predominantly aligned to a canonical frame where the ground plane corresponds to $z=0$. Following TRELLIS (Xiang et al., 2025), we encode each asset into occupancy grids and structured latents that are suitable for training the flow transformer. The structured latents can be decoded into high-quality triangle meshes or Gaussian splats using the SLat decoders.
To enable pose-grounded generation, we render 24 conditioning images from random viewpoints for each asset, with augmented focal lengths ranging from 24 mm to 200 mm. We then convert the camera pose and the coarse structure into a 3D occupancy grid and corresponding UV volume $\{\mathbf{x}_{i},\mathbf{u}_{i}(\bm{\theta})\}_{i=1}^{L}$, as described in the main paper. We first train a 3D VAE to encode the full UV volume of shape $64\times 64\times 64$ into 3D latents of shape $16\times 16\times 16$, which are concatenated with occupancy latents for training the first flow model $\mathbf{\mathcal{G}_{S}}$. Then, a second flow model $\mathbf{\mathcal{G}_{L}}$ is trained to generate the structured latents $\{\mathbf{f}_{i}\}_{i=1}^{L}$ based on the occupancy grid.
@@
We initialize our models using the pretrained weights of TRELLIS. During training, we apply classifier-free guidance (Ho & Salimans, 2022) (CFG) with a drop rate of 0.1. Both $\mathbf{\mathcal{G}_{S}}$ and $\mathbf{\mathcal{G}_{L}}$ are trained using AdamW (Loshchilov & Hutter, 2017) at a fixed learning rate of $1\times 10^{-4}$ for 500k and 100k steps, respectively. Training completes in approximately one week on 32 GPUs. At inference time, we use 25 sampling steps with classifier-free guidance strengths of 7.5 and 3.0 for $\mathbf{\mathcal{G}_{S}}$ and $\mathbf{\mathcal{G}_{L}}$, respectively.
@@
Figure 8: Qualitative comparison: Generative Reconstruction vs. 3D Generation. Unlike standard 3D generation, which aims to create novel objects from images, our generative reconstruction is specifically designed to accurately recreate a particular object that replicates the visible regions of the input image while maintaining diversity in the invisible regions. This difference in objectives is crucial, as prior 3D generators (Xiang et al., 2025) are not optimized for this task and often produce artifacts such as color drift and texture inconsistencies. As demonstrated in the canonical front and back views, our method’s pose-aligned image conditioning effectively reduces these issues, resulting in a reconstruction that remains true to the input.
@@ A.2 Occlusion-aware conditioning
We apply two modifications to both flow transformers to incorporate the mask. First, inspired by Amodal3R (Wu et al., 2025), we modulate the attention weight matrix during global condition injection via cross-attention using the mask. Specifically, the attention weight for each input token is computed by patching the mask to match the DINOv2 tokens and calculating the ratio of unmasked pixels in each patch. The logarithm of the weight values is added to the attention logits before applying the softmax operation. Second, for the geometry and appearance flow model that takes additional pose-aligned features, we concatenate the input image with the mask as an additional channel before feeding it into the convolution layer in the second stage.
@@
Reconstruction methods like LRM (Hong et al., 2023) typically generate a single 3D object from one image. However, our approach not only surpasses LRM in terms of reconstruction quality but also produces multiple plausible interpretations—what we refer to as generative reconstruction. Since standard metrics do not effectively capture this diversity, we offer qualitative comparisons in Figure 7. The results demonstrate that our framework creates diverse 3D models with visible regions that consistently align with the input image. In contrast, conventional 3D generation methods (Xiang et al., 2025) often struggle to maintain consistency in the visible regions of the objects they create. Furthermore, while directly processing an image with multiple objects may still yield a scene, the generation quality is often degraded as such inputs are out-of-distribution. In contrast, our method allows for component-aligned compositional reconstruction, as demonstrated by the additional examples in Figure 9.
A.4 Qualitative comparison with 3D generative methods
Since TRELLIS (Xiang et al., 2025) does not provide an explicit object-centric camera pose for the input image and post-hoc alignment is unreliable, we cannot conduct fair quantitative evaluation of reprojection consistency. We therefore focus on qualitative comparisons in Figure 8. For both methods, we render canonical object from front and back views. As shown, TRELLIS struggles to maintain texture consistency with the input image, exhibiting notable color drifting on the candle statue, teddy bear, and the lamb. In contrast, our method closely preserves the appearance of input. Importantly, although we incorporate pixel-level cues via back-projection, the model does not simply copy pixels: the generated back views remain coherent with the front, indicating learned view-consistent geometry and appearance. These results demonstrate that our mechanism effectively mitigates color drifting and texture inconsistency. We hypothesize that this limitation is common among 3D generators that lack localized, pixel-attended conditioning, including follow-up works of TRELLIS. We hope these findings can inspire future designs of 3D generators.
Appendix B Use of Large Language Models (LLMs)
We used LLMs to assist with writing. The authors drafted the manuscript, and LLMs were employed for grammar correction, copy-editing, and improving clarity and fluency. No novel scientific ideas, analyses, or experimental results were produced by LLMs. All LLM-assisted text was reviewed, revised, and approved by the authors, who take full responsibility for the content.

### v2

@@ 6 Discussion
We present Cupid, a generative reconstruction framework that unifies 3D generation and reconstruction by jointly modeling canonical 3D object and camera pose that formulate a given observation. By incorporating the missing camera pose prior and pose-aligned conditioning into 3D generators, this simple yet effective design enables Cupid to significantly improve geometric and photometric consistency with the input image for single-view reconstruction.
Despite strong performance and consistent gains over prior methods, Cupid has limitations. First, it requires object masks like existing 3D generation methods, and boundary errors in real images can degrade reconstruction quality. Second, lighting can be baked into the appearance, which can be improved with better material–light disentanglement. Third, our synthetic training images are mostly centered, making off-centered objects more challenging in real scenes. Nevertheless, these are not fundamental limitations and can be alleviated with better data and supervision.
Looking forward, this generative reconstruction framework enables promising test-time extensions, particularly for multi-view scenarios. From multiple images, our method refines 3D reconstructions to align with all observations by fusing a shared object latent during sampling, similar to Multi-Diffusion [bar2023multidiffusion]. As shown in Fig. 7, this yields an SfM-like system, though challenges like misaligned 3D orientations [lu2025orientation] from diverse views require advanced fusion schemes. Moreover, Cupid’s joint modeling enables bidirectional capabilities: flexible generation with known poses or pose estimation from images of given objects, facilitating mixed reality and embodied AI applications.
@@ Appendix A Appendix
Figure 8: Diversity of our Generative Reconstruction. We visualize canonical front and back views of generated 3D objects using random seeds (1-4). Given a single input image, our model synthesizes diverse hypotheses for unobserved regions while remaining highly consistent with visible regions. In contrast, the base 3D generator [xiang2025structured] struggles to conform to the input image, or produce less diverse unobserved regions.
@@ A.1 Implementation
We train our model using several datasets, including ABO [collins2022abo], HSSD [khanna2024habitat], 3D-FUTURE [fu20213d], and a subset of Objaverse-XL [deitke2023objaverse], totaling approximately 260K 3D assets. These artist-curated datasets are predominantly aligned to a canonical frame where the ground plane corresponds to $z=0$. Following TRELLIS [xiang2025structured], we encode each asset into occupancy grids and structured latents that are suitable for training the flow transformer. The structured latents can be decoded into high-quality triangle meshes or Gaussian splats using the SLat decoders. To enable diverse camera generation, we render 24 conditioning images from random viewpoints for each asset, with augmented focal lengths ranging from 24 mm to 200 mm.
Pose Encoding and Recovery.
Following the methodology outlined in the main text, we represent the camera pose and coarse scene geometry as a voxelized 3D occupancy grid paired with a UV coordinate volume, denoted as $\{\mathbf{x}_{i},\mathbf{u}_{i}(\bm{\theta})\}_{i=1}^{L}$. To facilitate efficient learning, we employ a 3D VAE to compress the input UV volume (resolution $64^{3}$) into a compact latent representation (resolution $16^{3}$). These pose latents are concatenated with occupancy embeddings to condition the training of the structural flow model, $\mathbf{\mathcal{G}_{S}}$. For pose recovery, the latents generated by $\mathbf{\mathcal{G}_{S}}$ are decoded back into the UV volume. We then estimate the camera extrinsics $\mathbf{R}|\mathbf{t}]$ and intrinsics $\mathbf{K}$ using the Perspective-n-Point (PnP) algorithm in Algo. 1. We evaluate the fidelity of this reconstruction in Tab. 5, reporting the mean and 95th percentile for Reprojection Error in normalized pixel coordinates, Relative Translation Error (RTE), Relative Rotation Error (RRE), and Relative Field-of-View (RFov) error in degrees. With the solved camera pose, we finally train the latent flow model $\mathbf{\mathcal{G}_{L}}$ to synthesize the structured features $\{\mathbf{f}_{i}\}_{i=1}^{L}$ conditioned on the occupancy and the pose-aligned visual features.
Table 5: Pose Reconstruction Fidelity. We evaluate the accuracy of camera poses recovered from the decoded UV volumes via the PnP algorithm. We report the Reprojection Error in normalized pixels, RRE, RTE, and RFov in degrees.
Metric Reproj. Err. RRE RTE RFov
Mean 0.0009 0.46 0.45 0.15
95% Quantile 0.0011 1.14 1.14 0.31
Algorithm 1 Perspective-n-Point solver via Direct Linear Transformation (DLT)
$\mathcal{X}$: Sets of 3D coordinates $\{\mathbf{x}_{i}\}_{i=1}^{N}$
$\mathcal{U}$: 2D coordinates $\{\mathbf{u}_{i}\}_{i=1}^{N}$
$N\leftarrow|\mathcal{X}|$ $\triangleright$ Number of correspondences
$\mathbf{A}\leftarrow\mathbf{0}_{2N\times 12}$ $\triangleright$ Initialize coefficient matrix
for $i=1$ to NN do
  $\tilde{\mathbf{x}}_{i}\leftarrow[\mathbf{x}_{i}^{\top},1]^{\top}$ $\triangleright$ Homogeneous coordinates
  $u_{i},v_{i}\leftarrow\mathbf{u}_{i}$
  $\mathbf{A}[2i-1,:]\leftarrow[\tilde{\mathbf{x}}_{i}^{\top},\mathbf{0}^{\top},-u_{i}\tilde{\mathbf{x}}_{i}^{\top}]$
  $\mathbf{A}[2i,:]\quad\leftarrow[\mathbf{0}^{\top},\tilde{\mathbf{x}}_{i}^{\top},-v_{i}\tilde{\mathbf{x}}_{i}^{\top}]$
end for
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}]\leftarrow\text{SVD}(\mathbf{A})$
$\mathbf{p}\leftarrow\mathbf{V}[:,-1]$ $\triangleright$ Solution is the last column of $\mathbf{V}$
$\mathbf{P}\leftarrow\text{reshape}(\mathbf{p},3,4)$ $\triangleright$ Camera matrix $\mathbf{P}\in\mathbb{R}^{3\times 4}$
if $\det(\mathbf{P}_{1:3,1:3})<0$ then $\triangleright$ Fix scale/sign ambiguity
  $\mathbf{P}\leftarrow-\mathbf{P}$
end if
$\mathbf{K},\mathbf{R},\mathbf{t}]\leftarrow\text{RQ}(\mathbf{P}) \triangleright$ Decompose camera matrix
return $\mathbf{K},\mathbf{R},\mathbf{t}$
Training Details.
@@
We initialize our models using the pretrained weights of TRELLIS. During training, we apply classifier-free guidance [ho2022classifier] (CFG) with a drop rate of 0.1. Both $\mathbf{\mathcal{G}_{S}}$ and $\mathbf{\mathcal{G}_{L}}$ are trained using AdamW [loshchilov2017decoupled] at a fixed learning rate of $1\times 10^{-4}$ for 500k and 100k steps, respectively. Training completes in approximately one week on 32 GPUs. At inference, we use 25 sampling steps with classifier-free guidance strengths of 7.5 and 3.0 for $\mathbf{\mathcal{G}_{S}}$ and $\mathbf{\mathcal{G}_{L}}$, respectively.
@@
Figure 9: Qualitative comparison: Generative Reconstruction vs. 3D Generation. Unlike standard 3D generation, which aims to create novel objects from images, our generative reconstruction is specifically designed to accurately recreate a particular object that replicates the visible regions of the input image while maintaining diversity in the invisible regions. This difference in objectives is crucial, as prior 3D generators [xiang2025structured] are not optimized for this task and often produce artifacts such as color drift and texture inconsistencies. As demonstrated in the canonical front and back views, our method’s pose-aligned image conditioning effectively reduces these issues, resulting in a reconstruction that remains true to the input.
@@ A.2 Occlusion-aware conditioning
We apply two modifications to both flow transformers to incorporate the mask. First, inspired by Amodal3R [wu2025amodal3r], we modulate the attention weight matrix during global condition injection via cross-attention using the mask. Specifically, the attention weight for each input token is computed by patching the mask to match the DINOv2 tokens and calculating the ratio of unmasked pixels in each patch. The logarithm of the weight values is added to the attention logits before applying the softmax operation. Second, for the geometry and appearance flow model that takes additional pose-aligned features, we concatenate the input image with the mask as an additional channel before feeding it into the convolution layer in the second stage.
@@
Reconstruction methods like LRM [hong2023lrm] typically generate a single 3D object from one image. However, our approach not only surpasses LRM in terms of reconstruction quality but also produces multiple plausible interpretations—what we refer to as generative reconstruction. Since standard metrics do not effectively capture this diversity, we offer qualitative comparisons in Figure 8. The results demonstrate that our framework creates diverse 3D models with visible regions that consistently align with the input image. In contrast, conventional 3D generation methods [xiang2025structured] often struggle to maintain consistency in the visible regions of the objects they create. Furthermore, while directly processing an image with multiple objects may still yield a scene, the generation quality is often degraded as such inputs are out-of-distribution. In contrast, our method allows for component-aligned compositional reconstruction, as demonstrated by the additional examples in Figure 10.
A.4 Qualitative comparison with 3D generators
Since TRELLIS [xiang2025structured] does not provide an explicit object-centric camera pose for the input image and post-hoc alignment is unreliable, we cannot conduct fair quantitative evaluation of reprojection consistency. We therefore focus on qualitative comparisons in Figure 9. For both methods, we render canonical object from front and back views. As shown, TRELLIS struggles to maintain texture consistency with the input image, exhibiting notable color drifting on the candle statue, teddy bear, and the lamb. In contrast, our method closely preserves the appearance of input. Importantly, although we incorporate pixel-level cues via back-projection, the model does not simply copy pixels: the generated back views remain coherent with the front, indicating learned view-consistent geometry and appearance. These results demonstrate that our mechanism effectively mitigates color drifting and texture inconsistency. We hypothesize that this limitation is common among 3D generators that lack localized, pixel-attended conditioning, including follow-up works of TRELLIS. We hope these findings can inspire future designs of 3D generators.